OpenStudy (anonymous):

evaluate 8C3 thr numbers are small and on the lower end of the C

OpenStudy (anonymous):

use a calc

OpenStudy (anonymous):

this is \[\frac{8\times 7\times 6}{3\times 2}=8\times 7=56\]

OpenStudy (anonymous):

means the number of ways you can choose 3 items from a set of 8. reads "eight choose 3"

OpenStudy (anonymous):

thats what i got thank you.